OpenStudy (anonymous):

What is the number of moles in 42 g of NaCl? A. 0.61 mol NaCl B. 1.8 mol NaCl C. 0.72 mol NaCl D. 1.2 mol NaCl

OpenStudy (anonymous):

@abb0t can you help?

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

@Preetha

OpenStudy (aaronq):

\(\sf \Huge moles=\dfrac{mass}{Molar~mass}\)

OpenStudy (anonymous):

Would it be D @aaronq

OpenStudy (aaronq):

whats the molar mass of NaCl?

OpenStudy (nechirwan):

C. 0.72 mol NaCl

OpenStudy (nechirwan):

moles=mass/Molar mass =42/58.5 =0.72 mol